OpenStudy (anonymous):

There is an order of 400ml of 0.50% of solution. You have in stock 16 oz. bottle of 0.75% and 16 oz. bottle of 0.35%. How much of each do you need to mix together. I know the answer is 150ml of .75% and 250ml of .35% but don't know how to get it.

OpenStudy (whpalmer4):

You need 400 ml of 0.5% solution. How much of the substance of interest is present in that?

OpenStudy (whpalmer4):

400 ml * 0.5% * 1/100% = 2 ml Now, if we use \(a\) to represent the quantity of \(0.75\%\) stock solution, and \(b\) to represent the quantity of \(0.35\%\) stock solution, the amount of the SOI contributed by \(a\) is \(0.75*0.01*a = 0.0075a\) and similarly for \(b\) it is \(0.35*0.01*b = 0.0035b\). Together, the two add up to contribute 2 ml to the final solution, so \[0.0075a + 0.0035b = 2\] Also, we know that we are going to make 400 ml of solution from \(a\) and \(b\), so \[a+b = 400\] That gives us two equations in two unknowns: \[0.0075a + 0.0035b = 2\]\[a+b=400\]To make this a bit easier, I would get rid of the decimals by multiplying the first equation by 1000: \[1000*0.0075a + 1000*0.0035b = 1000*2\]\[75a+35b = 2000\] Next, solve the other equation for one variable in terms of the other: \[a = b-400\] Now substitute the right hand side of that wherever you see the left hand side of that in the first equation: \[75a+35b=2000\]\[75(b-400) + 35b = 2000\] I'm confident you can solve that equation for \(b\). Then plug your value obtained for \(b\) into \(a = b-400\) and find \(a\).

OpenStudy (anonymous):

I will work on it and see what I get thanks

OpenStudy (whpalmer4):

uh don't follow my work exactly or you'll get it wrong :-) \[a+b = 400\]\[a = 400-b\]not \(a = b-400\)!

OpenStudy (whpalmer4):

Geez, I really didn't do well here, there's another mistake, too :-( multiply first equation by 10000, not 1000: \[75a+35b=20000\]

OpenStudy (anonymous):

reworking it still

OpenStudy (anonymous):

not getting correct answer. This is what I have so far. 75(400-b) +35b=20000 -75b+30000+35b=20,000 -40b=17,000 b=-425 (this can't be right -not sure what is wrong)

OpenStudy (whpalmer4):

\[75(400-b) + 35b = 20000\]\[30000-40b = 20000\]\[10000=40b\]

OpenStudy (anonymous):

got it subtrated 3,000 instead of using 30,000 thanks